# Eagle: Exploring The Design Space for Multimodal LLMs with Mixture of Encoders

### v1

@@ 2.4 Fusion strategy
Existing MLLM frameworks have proposed various fusion strategies to ensemble the encoders in Table 3, with the hope that their domain-specific strengths can be leveraged. In all cases, improvements in MLLM performance have been reported with the fusion of vision encoders. However, the roles of the fusion strategies as part of their MLLM architecture innovations, have not been decoupled and clearly studied under an “apples to apples” comparison. It is thus not entirely clear how much improvement is from the fusion strategies themselves versus the improved representations from various encoders.
@@
We notice that existing popular fusion strategies, despite their variations in designs, can be broadly represented by the following several categories: (1) Sequence Append: directly appending the visual tokens from different backbones as a longer sequence [20, 33, 55]; (2) Channel Concatenation: concatenating the visual tokens along the channel dimension without increasing the sequence length [49]; (3) LLaVA-HR: injecting high-resolution features into low-resolution vision encoders using mixture-of-resolution adapter [59]; (4) Mini-Gemini: using the CLIP tokens as the low resolution queries to cross-attend another high-resolution vision encoder in the co-located local windows [45]. (5) Deformable Attention: a new baseline we introduce on top of Mini-Gemini, where the vanilla window attention is replaced with deformable attention [86]. Fig. 2 gives a detailed illustration of these fusion strategies. To better study them, we choose “CLIP+ConvNeXt” and “CLIP+ConvNeXt+SAM” as the base multi-encoder combinations to perform comparisons.
@@
Our study in Table 4 shows that Channel Concatenation achieves the best average performance while maintaining better throughput compared to sequence append. The “injection-based” methods, such as LLaVA-HR, Mini-Gemini and Deformable Attention, are in general less competitive on TextVQA [69] and OCRBench [57], performing worse than using ConvNeXt alone as the vision encoder. A plausible explanation is that the CLIP features continue to play a dominant role in the visual tokens. Although sequence append shows comparable performance to channel concatenation, it faces the challenge to handle more vision encoders due to the increasing sequence length. Hence, we choose direct channel concatenation as our fusion strategy considering its performance, expandability, and efficiency.
Table 4: Comparison of different fusion methods for different vision experts. SA, $C$, LH, MG and DA denote the fusion strategies: Sequence Append, Channel Concatenation, LLaVA-HR [59], Mini-Gemini [45] and Deformable Attention, respectively. “Token (V)” denotes the number of vision feature tokens fed into LLM. “#Tokens/s” denotes the speed of the whole pipeline. “Avg” denotes average. Note that Channel Concatenation achieves both the best performance and better throughput.
Vision Encoder(s) Fusion Token (V) #Tokens/s  GQA  MME MMMU  OCR  SQA POPE TextVQA   Avg
CLIP 1024 47.9 64.6 1531 35.7 496 72.8 87.9 64.0 644.4
ConvNeXt 1024 46.8 63.7 1433 37.0 527 72.0 87.9 71.7 652.3
CLIP + ConvNeXt SA 2048 46.1 64.6 1482 35.3 536 72.5 88.3 71.8 657.2
$C$ 1024 47.3 64.0 1486 36.0 533 72.7 88.6 71.9 658.2
LH 1024 47.0 62.9 1488 36.8 521 72.6 88.9 68.7 652.1
MG 1024 45.2 64.9 1481 36.7 474 72.8 88.2 68.6 646.7
DA 1024 47.3 63.6 1497 35.3 459 71.9 88.0 69.1 640.9
CLIP + ConvNeXt + SAM SA 3072 40.3 64.9 1439 35.0 504 72.7 88.2 71.3 649.2
$C$ 1024 46.3 64.6 1497 35.2 558 71.7 87.9 72.2 660.4
@@ 2.5 Vison-language Pre-Alignment
As shown in Table 3, vision backbones pre-trained exclusively on vision tasks (e.g., detection, OCR, and segmentation) are less competitive compared to encoder pre-trained on vision language alignment. This is possibly due to representational inconsistencies when integrated with large language models. Additionally, when combining different encoders, there is a gap between these encoders, creating difficulties in the training process. To address this feature inconsistency, we propose a Pre-Alignment training stage that first aligns each individual vision encoder with the same large language model, fostering better synergy between visual and linguistic capabilities.
Fig. 3 depicts our pre-alignment strategy. Instead of training a projector to simultaneously align multiple vision experts as in LLaVA’s [54] original pre-training strategy, we first align the representation of each individual expert with a smaller language model (Vicuna-7B in practice) using next-token-prediction supervision. As shown in Fig. 3, with pre-alignment, the whole training process consists of three steps: 1) training each pre-trained vision expert with their own projector on SFT data, while keeping the language model frozen; 2) combining all the vision experts from the first step and training only the projector with image-text pairs data; 3) training the whole model on the SFT data.
@@
Table 5: Experiments to show the effectiveness of Pre-Align. “OCR” denotes OCRBench [57].“Avg” denotes average.
CLIP Vision Expert (X) Unfreeze Pre-Align  MME MMMU  OCR  SQA POPE TextVQA   Avg
CLIP-448 SAM-1024 1495 35.1 292 71.8 85.6 59.7 593
1504 36.9 482 73.5 87.7 63.2 641
1554 36.7 505 73.8 88.2 62.9 649
CLIP-448 ConvNext-1024 1499 35.1 416 72.3 86.7 67.1 629
1522 36.6 550 72.7 87.9 71.9 667
1538 36.3 556 71.9 87.9 71.6 667
CLIP-448 Pix2Struct-1024 1500 35.8 467 71.9 86.1 60.9 627
1529 36.2 535 73.1 87.1 64.5 651
1564 35.3 541 73.3 86.6 63.9 652
CLIP-448 EVA-02-L-1024 1484 35.8 305 72.0 86.0 59.9 597
1537 37.0 474 72.6 87.5 62.6 639
1586 37.3 495 73.1 88.9 64.2 653
@@ 2.6 Extension to multi-experts
With the optimized strategies and training recipes of incorporating individual vision experts, we consider the incorporation of even more vision experts to push the limit. To conduct the searching in a systematic and principled manner, we adopt a step-by-step greedy strategy to incorporate additional vision experts.
We consider the vision experts discussed in Section 2.3 for experiments. We mark CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct, and EVA-02-L as A, B, C, D, E, and F, respectively. A round-robin scheme, as shown in Table 6, is adopted. We first use the two top performing vision encoders, CLIP and ConvNeXt, as the basis and gradually add one more vision encoder each time. In each round, the best-performing vision encoder combination is retained for the next round.
Table 6: Results of vision expert selection process. A, B, C, D, E and F denote CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct and EVA-02-L models respectively. “OCR” denotes OCRBench [57]. “Avg” denotes average. Under the same number of encoders, the best scores of various combination variants are bold.
#Encoders Encoder Combination  GQA  MME MMMU  OCR  SQA POPE TextVQA   Avg
2 A + B 64.0 1486 36.0 533 72.7 88.6 71.9 658.3
3 A + B + C 64.6 1497 35.2 558 71.7 87.9 72.2 660.3
A + B + D 65.7 1506 35.0 509 72.6 87.9 70.1 653.6
A + B + E 65.5 1459 35.3 565 73.1 87.9 70.8 660.1
A + B + F 64.7 1506 35.9 562 73.2 88.3 72.1 665.3
4 A + B + F + C 65.9 1498 36.6 540 72.0 88.8 71.6 662.6
A + B + F + D 65.6 1521 35.8 508 73.0 89.0 70.0 657.5
A + B + F + E 65.3 1539 37.1 552 73.0 87.4 71.3 666.1
5 A + B + F + E + C 65.7 1511 37.6 548 72.7 88.6 71.6 666.5
A + B + F + E + D 65.6 1512 36.0 528 72.9 88.4 70.5 660.0
6 A + B + F + E + C + D 65.18 1504 35.3 534 71.9 87.6 70.86 656.3
Table 6 reveals several insights. Generally, introducing additional vision encoders enhances the performance. This indicates that the distinct advantages of different encoders can be preserved and utilized; for example, integrating the EVA-02 encoder improves metrics on the POPE benchmark. Although individual metrics may vary, the aggregated performance shows an upward trend, as evidenced by normalized average metrics, suggesting that the overall efficacy of the system is enhanced with more encoders. Also, Table 6 shows that the best combination of vision experts are CLIP, ConvNeXt, SAM, Pix2Struct, and EVA-02. We will use this recipe in our final model.
Table 7: The advanced training data used in the supervised fine-tuning stage.
Note Data Source
Total Number (1,809k) LLaVA-1.5 [52] (665k), DocVQA [61] (39k), synDog-EN [35] (50k), ChartQA [60] (28k), DVQA [32] (25k), AI2D [34] (15k), ShareGPT-4V [8] (100k), laion-GPT4V [1] (11k), LVIS-Instruct4V [74] (220k), LRV-Instruct [51] (150k), Geo170k [25] (120k), LLaVAR [84] (20k), Visual7W [87] (70k), Open-Hermes 2.5 [72] (1809k)
@@ 3 Experiments
After meticulously developing our recipes, we have formulated the following principals for our model, Eagle: (1) integrating more vision experts with optimized training recipe; (2) combining multiple vision experts through direct channel concatenation; (3) pre-training the vision experts separately via pre-alignment. In this section, to further show the advantages of the Eagle models, we add more training data and compare Eagle against the current state-of-the-art MLLMs on different tasks.
@@ 3.1 Implementation details
Models. We use Vicuna-v1.5-7B [15], Llama3-8B [3] and Vicuna-v1.5-13B [15] as the language model. For vision encoder, following the results in Sec. 2.6, we denote the model with four vision encoders: CLIP, ConvNeXt, Pix2Struct and EVA-02 as Eagle-X4 and the model with an additional SAM vision encoder as Eagle-X5.
Augmented training data. Considering the training cost, we only use 934K data during the investigation and ablations, as listed in Sec. 2.1. In this section, inspired by recent practices [73, 45], we introduce more datasets, including LVIS-INSTRUCT4V [74], LRV [51], Geo170K [25], LLaVAR [84], Visual7w [87], and OpenHermes-2.5 [72], resulting in a total of 1,809k data for the supervised fine-tuning stage. See Table 7 for more details on this data mixture. To better perform an apple-to-apple comparison with recent open-source MLLM projects, we also train Eagle-X5 with the same data as Cambrian-1 [73], including 2.5 million and 7 million data for pre-training and supervised fine-tuning.
@@ 3.2 Main results
Evaluation on visual question answering tasks. We compare Eagle model series across three Visual Question Answering (VQA) benchmarks, including GQA [30], VQAv2 [26] and VizWiz [27]. As shown in Table 8, Eagle-X5 achieves state-of-the-art performance on GQA and VQAv2, underscoring the advantages of incorporating additional vision experts.
Table 8: Results with the advanced training data recipe. SQAI denotes ScienceQA-IMG [65].
@@
Vicuna-7B & Qwen-7B LLaVA-1.5 [52] 1510 64.3 58.6 - - 85.9 66.8 62.0* 50.0 78.5* 297 58.2 -
LLaVA-NeXt [53] 1519 67.4 70.2 34.6 35.8 86.5 70.1 64.2* 57.6 80.0* 490 64.9 -
InternVL [13] 1525 - 65.4 - - 86.4 - 62.9* 52.5 79.3* - 57.0 -
LLaVA-HR [59] 1554 - 64.2 - - 87.6 65.1 64.2* 48.7 81.9* - 67.1 -
Monkey [47] - - - - - - - 60.7* 61.2* 80.3* 514 67.6 65.1
Mini-Gemini [45] 1523 65.8 - 32.2 36.8 - 71.1 64.5* - - 477 65.2 -
Eagle-X4 1561 67.8 73.4 37.3 34.9 88.4 70.4 64.8* 50.8 83.4* 540 70.9 67.5
+Pre-Align 1561 68.1 73.4 38.8 36.8 88.7 70.8 64.9* 51.0 83.7* 544 71.9 67.9
@@
+Pre-Align 1579 68.8 73.5 39.0 36.4 88.5 71.6 64.0* 56.0 83.6* 554 71.9 67.0
Vicuna-13B LLaVA-1.5 [52] 1531 67.7 61.6 - 36.4 85.9 71.6 63.3* 53.6 80.0* 331 61.3 -
LLaVA-NeXt [53] 1575 70.0 71.9 35.3 36.2 86.2 73.5 65.4* 60.5 82.8* 514 67.1 62.2
InternVL [13] 1546 - - - - 87.1 - 63.9* 54.6 80.2* 517 58.7 -
LLaVA-UHD [79] 1535 68.0 - - - 89.1 72.0 65.2* 56.1 81.7* - 67.7 -
LLaVA-HR [59] 1540 - 64.5 - - 87.8 68.1 64.8* 57.9 82.6* - 68.1 -
Mini-Gemini [45] 1565 68.6 70.6 37.0 37.3 - 71.9 65.8* - - 466 65.9 56.6
Eagle-X4 1627 69.9 74.4 37.6 36.9 87.7 73.1 66.3* 56.2 83.7* 561 73.9 70.5
+Pre-Align 1621 69.9 74.9 39.4 37.3 89.1 73.5 64.9* 55.9 84.5* 581 74.3 72.2
@@
Evaluation on OCR and chart understanding tasks. To evaluate the OCR, document and chart understanding capabilities of Eagle, we benchmark our model on OCRBench [57], TextVQA [69], and ChartQA [60]. As illustrated in Table 8, our model significantly surpasses competitors on TextVQA, benefiting from its high-resolution architecture and integration of different vision encoders. Notably, Eagle maintains a straightforward design, supporting up to 1024 tokens without requiring complex tile decomposition of images.
@@
Fig. 5 shows some examples of OCR and document understanding cases. With high-resolution adaptation and more vision experts, our model can identifies small text within images and accurately extracts information according to the users’ instructions. To better understand the benefits of introducing experts pre-trained on other vision tasks,we visualize the results of a model with only the ConvNeXt and CLIP vision encoders, compared to the results of Eagle-X5 in Fig. 4. With the full set of vision encoders, the model can successfully correct mistakes, showing that even when equipped with high-resolution vision encoders pre-trained on vision-language alignment, the model’s abilities can still be enhanced by integrating additional vision experts pre-trained on diverse vision tasks.
@@
Evaluation on multimodal benchmarks. We evaluate Eagle on seven benchmarks for MLLMs to demonstrate its capabilities from different perspectives, including MME [24], MMBench [56], SEED [41], MathVista [58], MMMU [81], ScienceQA [65], and POPE [46]. Specifically, MME, MMBench, and SEED assess the overall performance on various real-world tasks based on reasoning, recognition, knowledge, and OCR. MMMU focuses on challenging problems from diverse domains that require college-level knowledge. POPE evaluates the visual hallucinations of MLLMs. The metrics used in our paper adhere to the default settings of these benchmarks. We report the perception score for MME, the en_dev split for MMBench, the image split of SEED, the test-mini split of MathVista, the val split of MMMU, the F1-score of POPE, and the image of SQA to align with the reported scores from other models.
@@
From the data presented in Table 8, Eagle consistently surpasses existing models across various MLLMs on SEED and MME, demonstrating the comprehensive knowledge and reasoning abilities of Eagle. With the help of vision encoders on object-centric tasks, Eagle also achieves the best performance on the POPE benchmark. Additionally, the Pre-Alignment strategy discussed in Sec. 2.5 has been found to further enhance performance when integrating multiple task-specific vision backbones. This approach not only mitigates the inherent biases of each vision expert and the synergy between different modalities but also establishes a robust framework for the multiple-expert fusion.
@@
Comparison with Cambrian-1. Using the same pre-training and supervised fine-tuning datasets from Cambrian-1 [73], Eagle demonstrates superior performance across all the evaluated benchmarks without bells and whistles. As shown in Table 9, Eagle outperforms the Cambrian-1 counterparts considerably on the OCR and Chart category. Consistent improvements are also observed on the General, Knowledge and Vision-Centric categories, showing the robustness and generalization ability of the improved perception design in Eagle.
@@
Table 9: Results using Cambrian-1 training data. The Eagle-X5 models are trained using the same pre-training and supervised fine-tuning data as Cambrian-1 [73]. SQAI denotes ScienceQA-IMG [65]. RWQA denotes RealworldQA [78].
@@
Eagle 68.6 85.5 53.2 57.9 79.1 76.3 1677 81.0 75.6 64.9 75.4 77.2 62.4 78.8 83.0 68.3 67.0 69.5

### v2

@@ 2.4 Fusion strategy
Existing MLLM frameworks have proposed various mixture-of-vision-encoder strategies, with the hope that their domain-specific strengths can be leveraged. In all cases, improvements in MLLM performance have been reported with the fusion of vision encoders. However, the roles of the fusion strategies as part of their MLLM architecture innovations, have not been decoupled and clearly studied under an “apples to apples” comparison. It is thus not entirely clear how much improvement is from the fusion strategies themselves versus the improved representations from various encoders.
@@
We notice that existing popular fusion strategies, despite their variations in designs, can be broadly represented by the following several categories: (1) Sequence Append: directly appending the visual tokens from different backbones as a longer sequence (Fan et al., 2024; Kar et al., 2024); (2) Channel Concatenation: concatenating the visual tokens along the channel dimension without increasing the sequence length (Lin et al., 2023b; Karamcheti et al., 2024); (3) LLaVA-HR: injecting high-resolution features into low-resolution vision encoders using mixture-of-resolution adapter (Luo et al., 2024); (4) Mini-Gemini: using the CLIP tokens as the low-resolution queries to cross-attend another high-resolution vision encoder in the co-located local windows (Li et al., 2024b). (5) Deformable Attention: a new baseline we introduce on top of Mini-Gemini, where the vanilla window attention is replaced with deformable attention (Zhu et al., 2021). Fig. 2 gives a detailed illustration of these fusion strategies. To better study them, we choose “CLIP+ConvNeXt” and “CLIP+ConvNeXt+SAM” as the base multi-encoder combinations to perform comparisons.
Table 4: Comparison of different fusion methods for different vision experts. “#Token(V)” denotes the number of visual tokens. “#Tokens/s” denotes the inference throughtput of the whole pipeline.
Vision Encoders Fusion #Token(V) #Tokens/s #Params Avg.
CLIP + ConvNeXt Seq. Append 2048 46.1 1200M 690.5
Channel Concat. 1024 47.3 1184M 681.5
LLaVA-HR 1024 47.0 1219M 678.7
Mini-Gemini 1024 45.3 1201M 672.5
Deformable Attn. 1024 47.3 1201M 674.3
CLIP + ConvNeXt + SAM Seq. Append 3072 40.3 1529M 686.2
Channel Concat. 1024 46.3 1495M 690.4
@@
Our study in Table 4 shows that Channel Concatenation stands out with the best performance, expandability, and efficiency. The “injection-based” methods, such as LLaVA-HR, Mini-Gemini and Deformable Attention, are in general less competitive on TextVQA (Singh et al., 2019) and OCRBench (Liu et al., 2023f), performing worse than using ConvNeXt alone as the vision encoder. Although sequence append shows comparable performance to channel concatenation, it faces the challenge of handling significantly increased sequence lengths with additional vision encoders.
@@ 2.5 Vison-language Pre-Alignment
As shown in Table 3, encoders pre-trained exclusively on vision tasks (e.g., detection, OCR, and segmentation) are less competitive compared to those pre-trained on vision language alignment. This is possibly due to representational inconsistencies when integrated with large language models. Additionally, when combining different encoders, there is a gap between these encoders, creating difficulties in the training process. To address this feature inconsistency, we propose a Pre-Alignment training stage that first aligns each individual vision encoder with the same large language model, fostering better synergy between visual and linguistic capabilities.
Fig. 3 depicts our pre-alignment strategy. Instead of training a projector to simultaneously align multiple vision experts as in LLaVA’s (Liu et al., 2023c) original pre-training strategy, we first align the representation of each individual expert with a smaller language model (Vicuna-7B in practice) using next-token-prediction supervision. As shown in Fig. 3, with pre-alignment, the whole training process consists of three steps: 1) training each pre-trained vision expert with their own projector, while keeping the language model frozen; 2) combining all vision experts from the first step and training both the projector and vision experts; 3) training the whole model on SFT data.
Table 5: The effectiveness of Pre-alignment.
CLIP Vision Expert (X) Unfreeze Pre-align Avg.
CLIP-448 SAM-1024 630.6
648.5
662.5
672.3
CLIP-448 ConvNext-1024 652.0
670.1
681.5
686.2
CLIP-448 Pix2Struct-1024 653.5
665.7
673.7
680.4
CLIP-448 EVA-02-L-1024 630.2
645.2
659.2
668.2
@@ 2.6 Extension to multi-experts
With the optimized strategies and training recipes of incorporating individual vision experts, we consider the incorporation of even more vision experts to push the limit. To conduct the search in a systematic manner, we adopt a step-by-step greedy strategy to incorporate additional vision experts.
We consider the vision experts discussed in Section 2.3 for experiments. We mark CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct, and EVA-02-L as CL, CN, SA, DI, PS, and EV, respectively. A round-robin scheme, as shown in Fig. 4, is adopted. We first use the two top-performing vision encoders, CLIP and ConvNeXt, as the basis and gradually add one more vision encoder each time. In each round, the best-performing vision encoder combination is retained for the next round.
#Encoder Encoder Combination Config #Params (M) FLOPs (G) Img/Sec Avg.
2 CL + CN X2 1155.2 3347.2 18.1 681.5
3 CL + CN + DI 1460.6 3659.9 15.1 685.4
CL + CN + SA 1463.9 4657.8 8.8 690.4
CL + CN + PS 1669.6 4373.2 6.9 685.1
CL + CN + EV X3 1459.6 4280.9 9.1 690.7
4 CL + CN + EV + DI 1765.1 4593.6 8.3 688.0
CL + CN + EV + SA 1768.4 5591.5 5.9 689.4
CL + CN + EV + PS X4 1974.1 5306.9 5.0 694.6
5 CL + CN + EV + PS + DI 2279.5 5619.5 4.7 684.7
CL + CN + EV + PS + SA X5 2282.8 6617.4 3.8 697.1
6 CL + CN + EV + PS + SA + DI X6 2588.2 6930.1 3.6 686.8
Figure 4: Results of vision expert selection process. CL, CN, EV, PS, SA and DI denote CLIP, ConvNeXt, EVA-02, Pix2Struct, SAM and DINOv2, respectively. (Left) The performance of various vision encoder combinations is presented, highlighting how different pairings influence overall effectiveness. “#Parames”, “FLOPs” and “Img/Sec” denote the model size, complexity and throughput (bs=4) of the vision encoder. (Right) The curve illustrates the average score as the number of vision encoders increases. Each point on the curve represents the best-performing combination for the corresponding number of vision encoders. No Pre-Alignment is used in this comparison.
Fig. 4 reveals several insights. Generally, introducing additional vision encoders enhances the performance. This indicates that the distinct advantages of different encoders can be preserved and utilized; for example, integrating the EVA-02 encoder improves metrics on the POPE benchmark. Although individual metrics may vary, the aggregated performance shows an upward trend, as evidenced by normalized average metrics, suggesting that the overall efficacy of the system is enhanced with more encoders. Also, Fig. 4 shows that the best combination of vision experts are CLIP, ConvNeXt, SAM, Pix2Struct, and EVA-02. We will use this recipe in the final model.
@@ 3 Experiments
In this section, we take the findings and the best-explored design from Section 2 and compare them against the current state-of-the-art MLLMs on different tasks.
@@ 3.1 Implementation details
Language models. We use Vicuna-v1.5-7B (Chiang et al., 2023), Llama3-8B (AI@Meta, 2024) and Vicuna-v1.5-13B (Chiang et al., 2023) as the language models.
Vision encoders. We follow the best X4 and X5 configurations, where the interpolated CLIP-448 and pre-aligned vision experts are channel-concatenated, and trained following the exact best training recipes in Figure 3 and Table 5.
Training recipe. On Eagle1.8M, we follow the base recipe in Section 2.1 with encoder learning rate the same as SFT (2e-5). On Cambrian data, we follow Tong et al. (2024) with PT/SFT bs=1024.
@@ 3.2 Main results
Evaluation on visual question answering tasks. We compare Eagle model series across three Visual Question Answering (VQA) benchmarks, including GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017) and VizWiz (Gurari et al., 2018). As shown in Table 6, Eagle-X5 achieves state-of-the-art performance on GQA and VQAv2, underscoring the advantages with additional vision experts.
Table 6: Main results with base training data. SQAI denotes image split of ScienceQA.
@@
Vicuna-7B & Qwen-7B LLaVA-1.5 (Liu et al., 2023c) 1510 64.3 58.6 - - 85.9 66.8 62.0* 50.0 78.5* 297 58.2 -
LLaVA-NeXt (Liu et al., 2024a) 1519 67.4 70.2 34.6 35.8 86.5 70.1 64.2* 57.6 80.0* 490 64.9 -
InternVL (Chen et al., 2023f) 1525 - 65.4 - - 86.4 - 62.9* 52.5 79.3* - 57.0 -
LLaVA-HR (Luo et al., 2024) 1554 - 64.2 - - 87.6 65.1 64.2* 48.7 81.9* - 67.1 -
Monkey (Li et al., 2024c) - - - - - - - 60.7* 61.2* 80.3* 514 67.6 65.1
Mini-Gemini (Li et al., 2024b) 1523 65.8 - 32.2 36.8 - 71.1 64.5* - - 477 65.2 -
@@
Eagle-X5 (+Pre-Align) 1582 69.7 73.7 38.2 38.0 88.7 71.9 64.6* 58.7 83.6* 566 71.9 69.3
Vicuna-13B LLaVA-1.5 (Liu et al., 2023c) 1531 67.7 61.6 - 36.4 85.9 71.6 63.3* 53.6 80.0* 331 61.3 -
LLaVA-NeXt (Liu et al., 2024a) 1575 70.0 71.9 35.3 36.2 86.2 73.5 65.4* 60.5 82.8* 514 67.1 62.2
InternVL (Chen et al., 2023f) 1546 - - - - 87.1 - 63.9* 54.6 80.2* 517 58.7 -
LLaVA-UHD (Xu et al., 2024) 1535 68.0 - - - 89.1 72.0 65.2* 56.1 81.7* - 67.7 -
LLaVA-HR (Luo et al., 2024) 1540 - 64.5 - - 87.8 68.1 64.8* 57.9 82.6* - 68.1 -
Mini-Gemini (Li et al., 2024b) 1565 68.6 70.6 37.0 37.3 - 71.9 65.8* - - 466 65.9 56.6
@@
Evaluation on OCR and chart understanding tasks. To evaluate the OCR, document, and chart understanding capabilities of Eagle, we benchmark our model on OCRBench (Liu et al., 2023f), TextVQA (Singh et al., 2019), and ChartQA (Masry et al., 2022). As illustrated in Table 6, our model significantly surpasses competitors on TextVQA, benefiting from its high-resolution architecture and integration of different vision encoders. Notably, Eagle maintains a straightforward design, supporting up to 1024x1024 resolution without requiring complex tile decomposition of images.
@@
Fig. 5 shows some examples of OCR and document understanding cases. With high-resolution adaptation and more vision experts, our model can identify small text within images and accurately extract information according to the users’ instructions. To better understand the benefits of introducing experts pre-trained on other vision tasks, we visualize the results of a model with only the ConvNeXt and CLIP vision encoders, compared to the results of Eagle-X5 in Fig. 5. With the full set of vision encoders, the model can successfully correct mistakes, showing that even when equipped with high-resolution vision encoders pre-trained on vision-language alignment, the model’s abilities can still be enhanced by integrating additional vision experts pre-trained on diverse vision tasks.
@@
Evaluation on multimodal benchmarks. We evaluate Eagle on seven benchmarks for MLLMs to demonstrate its capabilities from different perspectives, including MME (Fu et al., 2023), MMBench (Liu et al., 2023e), SEED (Li et al., 2023c), MathVista (Lu et al., 2024), MMMU (Yue et al., 2024), ScienceQA (Saikh et al., 2022), and POPE (Li et al., 2023e). Specifically, MME, MMBench, and SEED assess the overall performance on various real-world tasks based on reasoning, recognition, knowledge, and OCR. MMMU focuses on challenging problems from diverse domains that require college-level knowledge. POPE evaluates the visual hallucinations of MLLMs. The metrics used in our paper adhere to the default settings of these benchmarks. We report the perception score for MME, the en_dev split for MMBench, the image split of SEED, the test-mini split of MathVista, the val split of MMMU, the F1-score of POPE, and the image split of SQA to align with the reported scores from other models.
Table 7: Comparison between different training strategies. “1 epoch” means we train Eagle for 1 epoch in the supervised fine-tuning stage. ‘unlock*” means we unlock vision encoders in the pre-training stage. More details in Table. 13.
Config Summary Pre-align Pre-train Fine-tune Avg.
1 epoch LLaVA-595K Eagle1.8M 697.1
2 epoch LLaVA-595K Eagle1.8M 698.3
1 epoch, unlock* LLaVA-595K Eagle1.8M 698.0
1 epoch, unlock* LLaVA-595K+Eagle1.8M Eagle1.8M 699.5
1 epoch Eagle1.8M LLaVA-595K Eagle1.8M 706.6
1 epoch, unlock* Eagle1.8M LLaVA-595K Eagle1.8M 707.1
1 epoch, unlock* LLaVA-595K+Eagle1.8M LLaVA-595K Eagle1.8M 707.8
1 epoch, unlock* LLaVA-595K+Eagle1.8M LLaVA-595K+Eagle1.8M Eagle1.8M 708.9
@@
From the data presented in Table 6, Eagle consistently surpasses existing models across various MLLMs on SEED and MME, demonstrating the comprehensive knowledge and reasoning abilities of Eagle. With the help of vision encoders on object-centric tasks, Eagle also achieves the best performance on the POPE benchmark. Additionally, the Pre-Alignment strategy discussed in Sec. 2.5 has been found to further enhance performance when integrating multiple task-specific vision backbones. This approach not only mitigates the inherent biases of each vision expert and the synergy between different modalities but also establishes a robust framework for multiple-expert fusion.
Study on more advanced training recipes. Table. 7 presents our step-by-step experiments to study the training recipes. We found that the best recipe is to first pre-align each vision expert on LLaVA-595K + Eagle1.8M. In the pretraining stage, we combine all vision experts from the first step and training both the projector and vision experts on LLaVA-595K + Eagle1.8M. Finally, we train the whole model on the Eagle1.8M.
@@
Comparison with Cambrian-1. Using the same pre-training and supervised fine-tuning datasets from Cambrian-1 (Tong et al., 2024), Eagle demonstrates superior performance across all the evaluated benchmarks without bells and whistles. As shown in Table 8, Eagle outperforms the Cambrian-1 counterparts considerably for the OCR and Chart category. Consistent improvements are also observed for the General, Knowledge, and Vision-Centric categories, showing the robustness and generalization ability of the improved perception design in Eagle.
@@
Table 8: Results using the same training data as Cambrian-1 (Tong et al., 2024). SQAI denotes ScienceQA-IMG (Saikh et al., 2022). RWQA denotes the RealworldQA (xAI, 2024).
@@
Eagle-X5 68.6 85.5 53.2 57.9 79.1 76.3 1677 81.0 75.6 64.9 75.4 77.2 62.4 78.8 83.0 59.8 50.0 69.5
4 Related work
Multimodal large language models. Our work is related to the general architecture design of multimodal large language models. Besides the line of representative open-source research mentioned in the introduction section, other notable families of MLLMs include, but are not limited to MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023a), Lynx (Zeng et al., 2023), Otter (Li et al., 2023b; a), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023b; Hong et al., 2024), VILA (Lin et al., 2023a), GPT-4V (Achiam et al., 2023), Gemini (Team et al., 2023), and Llama 3.1 (Dubey et al., 2024). Depending on how vision signals are integrated into the language model, MLLMs can be broadly categorized into “cross-modal attention” ones and “prefix-tuning” ones (Yin et al., 2024). The former injects visual information into different layers of LLMs using cross-modal attention (Alayrac et al., 2022; Li et al., 2023b), whereas the latter views the visual tokens as part of the language token sequence and directly append them with text embeddings (Liu et al., 2023d; Li et al., 2022; 2023d). Our model belongs to the prefix-tuning family by following a LLaVA-styled multimodal architecture. Considering that MLLM is a fast-growing field, readers are recommended to refer to more detailed studies and surveys such as Yin et al. (2024); Li et al. (2024a); Fei et al. (2024).
Vision encoder design for MLLMs. Our work is closely related to research focusing on improved vision encoder designs for MLLM. Early works (Liu et al., 2023d; Li et al., 2022) usually adopted vision encoder pre-trained on vision-language alignment tasks such as CLIP (Radford et al., 2021) and EVA-CLIP (Sun et al., 2023). Stronger vision encoders such as SigLIP (Zhai et al., 2023) and InternVL (Chen et al., 2023f) have been proposed to improve vision-language tasks with improved designs, larger model sizes, and better training recipes. Considering that the models are often pre-trained on low-resolution images and lack the ability to encode fine-grained details, higher resolution adaptation is often performed to increase the MLLM input resolution (Chen et al., 2023c; d; Beyer et al., 2024; Chen et al., 2024).
Besides higher resolution adaptation, models such as LLaVA-NeXT (Liu et al., 2024a), LLaVA-UHD (Xu et al., 2024), Monkey (Li et al., 2024c), InternLM-XComposer (Dong et al., 2024), and InternVL (Chen et al., 2024) use adaptive tiling to divide high-resolution input into lower resolution patches that are processed separately. Although the ability to handle higher resolution is similarly made possible with the introduction of additional vision experts, its spirit is slightly orthogonal to the tiling techniques in terms of diversifying the visual representations. Both techniques are compatible and can be combined together.
Our work is most related to existing models using multiple vision encoders for improved perception. Mini-Gemini (Li et al., 2024b) and LLaVA-HR (Luo et al., 2024) propose to fuse high-resolution visual features into the low-resolution visual tokens. Apart from the resolution issue, these pre-trained vision encoders may lack specific abilities such as reading text and localizing objects. Hence, a series of works have integrated vision models pre-trained on different tasks for more comprehensive capabilities. For example, Mousi (Fan et al., 2024), Prismatic VLM (Karamcheti et al., 2024) and Brave (Kar et al., 2024) fuse visual tokens from different vision encoders by concatenating along the channel or token direction. There are also more complex approaches, including knowledge distillation Ranzinger et al. (2024), augmenting the input prompt with the information output by the vision experts (Lee et al., 2024; He et al., 2024; Liu et al., 2024b), or using a routing network to assign input to proper vision experts (Zong et al., 2024). In particular, Prismatic VLM (Karamcheti et al., 2024) systematically explores the design space of MLLMs across various dimensions, including data, training recipes, and notably. It also includes the vision encoder ensemble as part of its enhancement strategies. However, it lacks a comprehensive ablation study and discussion addressing the challenges associated with combining multiple vision encoders.